# Hotfixing Large Language Models for Code

### v1

@@ 1. Introduction
Despite their strong performance, LLM4Code are usually found to exhibit undesired behaviors. For example, Jesse et al. (Jesse et al., 2023) find that LLM4Code can memorize and produce many known simple, stupid bugs (i.e., bugs that can be fixed with single-line changes, also known as SStuBs (Karampatsis and Sutton, 2020)). As another example, researchers have found that LLM4Code can generate a large number of privacy-sensitive information (Yang et al., 2024b), like email addresses and API keys. Considering the large population of users of LLM4Code, [^1] Reports show that GitHub Copilot has already been used by over 1.2 million developers. we hope to update the model as soon as possible to minimize the potential negative impacts on more users caused by these undesired behaviors. One intuitive way is to retrain the model on the updated dataset, by adding the fixed code or replacing the sensitive information with placeholders in the training data, which however, is extremely time-consuming and computationally expensive.
@@
This practical requirement motivates us to propose the concept of “hotfixing” LLM4Code. Hotfix, in the context of traditional software engineering, is an unplanned improvement to remediate the unwanted symptoms of the critical issue quickly (Hanna et al., 2024). Hotfix emphasizes less on correctness under all conditions but more on the time required to generate a plausible patch that hides the critical symptom without breaking the system. In the context of LLM4Code, retraining the model thoroughly can be quite time-consuming, we hope to first hotfix the model to mitigate the undesired behaviors and thus leave time for a complete fix (retraining). A good hotfix for LLM4Code should satisfy the following properties: (1) efficiency: the time and effort to obtain the hotfix should be significantly faster than retraining the model; (2) effectiveness: the hotfix should effectively mitigate the behavior while causing minimal impact to the functional correctness; and (3) easy deployment: the hotfix should minimize disruption and complexity during deployment.
@@
This paper implements hotfixing to mitigate an important type of undesired behavior: LLM4Code can complete many known buggy code; we aim at making LLM4Code not complete buggy code but complete the corresponding fixed code. First, we replicate the experiments of Jesse et al. (Jesse et al., 2023), confirming that models from a popular family (CodeGen) can memorize and produce many bugs, which is treated as undesired behaviors that should be mitigated. These known bugs have been fixed in open-source software; thus we hotfix the models and steer them to generate the corresponding fixed versions. To achieve this goal, we construct a dataset consisting of pairs of two code snippets: one containing bugs and the other with corresponding bugs being fixed. Then, we compute the diff (to get the deleted buggy code and the added fixed code) of each pair.
Considering the effectiveness requirement, the model after hotfixing is expected to (1) produce more fixed code, (2) produce less buggy code, and (3) retain knowledge about other code from the original model. We design different loss functions for the three objectives. For objective (1), we minimize the log-likelihood loss on the added code. For objective (2), we design a loss that can penalize the generation of the buggy code. For objective (3), inspired by the knowledge distillation studies (Tang et al., 2024; Shi et al., 2024), we minimize the KL divergence (Kullback and Leibler, 1951) of the token distribution of the hotfixed model and the original model. PEFT methods are designed to update the model with a small number of parameters, resulting in efficiency gains and making the deployment easier. So we employ four parameter-efficient fine-tuning (PEFT) methods for the efficiency and ease of deployment of hotfixing.
@@
We investigate a family of popular LLM4Code, CodeGen (Nijkamp et al., 2023), which is widely adopted in multiple studies (Jesse et al., 2023; Yang et al., 2024b; Li et al., 2023; Allal et al., 2023). By conducting experiments on models of two sizes (CodeGen-350M and CodeGen-2B), we gain the following insights about hotfixing LLM4Code. The optimal hotfixing setting (combining all three objectives and using LoRA for fine-tuning) can increase the generation of fixed code by up to 108.42% and decrease the generation of buggy code by up to 50.47%, which is more effective than the mitigation strategy adopted by Jesse et al. (Jesse et al., 2023). Although hotfixing reduces the model’s functional correctness on the HumanEval benchmark, statistical tests show that the difference is not statistically significant ($p$-values are greater than 0.05).
@@
  • Effectiveness: We implement a hotfix strategy that can effectively mitigate two types of undesired behaviors with no significant impact on the functional performance.
@@ 3.3. Learning Objectives
However, in the context of hotfixing LLM4Code, directly training models on the entire updated examples may not be the appropriate option. In hotfixing, not each token in the entire updated examples is equally important, and the model should pay more attention to the added fixed code. Besides, the model should also learn not to generate the original buggy code and the model should cause minimal impact on the outputs that are not relevant to the bug-fixing updates. We explain the three objectives and the corresponding loss function design as follows.
@@ 3.3.2. Learning to Avoid Undesired Code
In the above equation, $w_{t}^{-}$ is the weight in computing the loss on token $x_{t}$, where $w_{t}^{-}=1$ if $x_{t}$ is the desired code and $w_{t}^{-}=0$ otherwise. Unlike the desired code, we want to increase the model loss on the undesired code to avoid them from being generated. However, simply maximizing the loss on the undesired code is ineffective. Our experiment shows that maximizing this loss will make the model lose its ability in generating meaningful code (the hotfixed models generate zero buggy code and zero fixed code), which is not desired in hotfixing. Instead, we propose to minimize the new loss function: $\frac{L_{guided}}{L_{guided}+L_{unlearn}}$. This loss function guides the optimizer to find a balance between minimizing the loss on desired code $L_{guided}$ and maximizing the loss on undesirable code $L_{unlearn}$. Ideally, to minimize this loss, the optimizer works to reduce $L_{guided}$ which encourages fixed code generation and to increase $L_{unlearn}$ which discourages the generation of buggy code. To make the model learn the desired code and avoid the undesired code, we define the Dual loss as follows:
@@ 3.3.3. Learning to Retrain Knowledge of Other Code
The fact that tuning language models on a new dataset can lead to performance degradation on other aspects has been observed in previous studies. For example, Liu et al. (Liu et al., 2016) point out that optimizing models on a new dataset can improve the average performance (e.g., BLEU score) but forgo coherence and fluency. This degeneration is often diagnosed as an effect of deviating too much from the original pre-trained model during optimization. Consequently, we want the model to behave similarly to the original model on the code that is not relevant to the undesired behavior. This is similar to the target of knowledge distillation (Tang et al., 2024; Shi et al., 2024), which aims to train a student model to mimic the behavior of the teacher model. In knowledge distillation, Kullback-Leibler (KL) divergence (Kullback and Leibler, 1951) is widely used to measure the behavioral difference between the two models and the student model learns to minimize the difference (Shi et al., 2023; Kim and Rush, 2016; Kim et al., 2024). So we define a new loss function as follows:
@@ 7.2. Efficient-tuning for Models of Code
Weyssow et al. (Weyssow et al., 2024) use 5 parameter-efficient methods to fine-tune models for code generation tasks. We reuse their replication package to implement the first hotixing for mitigating undesired behavior. Prefix-tuning falls into the category of efficient-tuning methods for pre-trained models, which aim to utilize computationally-friendly techniques to efficiently adapt a model to a new dataset or task. Typical methods include in-context learning, prefix-tuning, adapter-tuning. Here we briefly introduce the related work about applying these methods to models of code. In-context learning (ICL) presents a small number of examples in the prompt to guide language models to conduct specific tasks. Gao et al. (Gao et al., 2023) evaluate ICL on code summarization, bug fixing, and program synthesis. They show that strategically designed context can lead to substantial improvements over widely-used demonstration construction methods. Huang et al. (Huang et al., 2023a) show that ICL can mitigate the social bias of code generated by language models. Wang et al. (Wang et al., 2022) apply prompt-tuning on three code intelligence tasks and show that prompt-tuning consistently outperforms fine-tuning in three investigated tasks. Wang et al. (Wang et al., 2023) evaluate adapter tuning for code search and summarization, showing that adapter tuning significantly outperforms full-model fine-tuning and effectively overcomes catastrophic forgetting. Choi and Lee (Choi and Lee, 2023) propose CodePrompt, using task-agnostic prefix tuning for program and language generation. He and Vechev (He and Vechev, 2023) use contrastive prefix-tuning (Qian et al., 2022) to control the security-level of code generators.
@@ 8. Conclusion and Future Work
In this study, we present the first investigation of hotfixing, a lightweight method to mitigate the undesired behavior of code generation models. We define three learning objectives in hotfixing and design multiple loss functions for each objective: (1) learn the desired behaviors, (2) unlearn the undesired behaviors, and (3) retain knowledge of other code. We evaluate four different PEFT techniques for hotfixing the models. We specifically focus on two types of undesired behavior that are prevalent in code generation models: (1) generating buggy code and (2) leaking privacy information. We conduct our experiments on the models in the CodeGen family. The empirical results demonstrate the effectiveness of hotfixing: the best combination leads to a significant increase in the generation of fixed code (up to 108.42%) and a notable reduction in the production of buggy code (up to 50.47%). We also conduct statistical analysis to confirm that hotfixing does not adversely impact the functional performance of the models on the HumanEval benchmark. We discuss the potential applications of the proposed method in various tasks, as well as its advantage in efficient deployment.

### v2

@@ 1. Introduction
Despite their strong performance, LLM4Code are usually found to exhibit undesired behaviors. For example, Jesse et al. (Jesse et al., 2023) find that LLM4Code can memorize and produce many known simple, stupid bugs (i.e., bugs that can be fixed with single-line changes, also known as SStuBs (Karampatsis and Sutton, 2020)). As another example, researchers have found that LLM4Code can generate a large number of privacy-sensitive information (Yang et al., 2024b), like email addresses and API keys. Considering the large population of users of LLM4Code, [^1] Reports show that GitHub Copilot has already been used by over 1.2 million developers. there exists a key requirement to update the model as soon as possible to minimize the potential negative impacts on more users caused by these undesired behaviors. One intuitive way is to retrain the model on the updated dataset, by adding the fixed code or replacing the sensitive information with placeholders in the training data, which however, is extremely time-consuming and computationally expensive.
@@
This practical requirement motivates us to explore the concept of “hotfixing” LLM4Code. Hotfix, in the context of traditional software engineering, is an unplanned improvement to remediate the unwanted symptoms of the critical issue quickly (Hanna et al., 2024). Hotfix emphasizes less on correctness under all conditions but more on the time required to generate a plausible patch that hides the critical symptom without breaking the system. In the context of LLM4Code, retraining the model thoroughly can be quite time-consuming, we hope to first hotfix the model to mitigate the undesired behaviors and thus leave time for a complete fix (retraining). A good hotfix for LLM4Code should satisfy the following properties: (1) efficiency: the time and effort to obtain the hotfix should be significantly faster than retraining the model; (2) effectiveness: the hotfix should effectively mitigate the behavior while causing minimal impact to the functional correctness; and (3) easy deployment: the hotfix should minimize disruption and complexity during deployment.
@@
He and Vechev (He and Vechev, 2023) propose the pioneering study on updating LLM4Code to reduce the generation of insecure code. They define three targets for such updates: the updated model should (1) produce more secure code, (2) produce less insecure code, and (3) remain the same for other code. To achieve these targets, they design a combination of three loss functions to train contrastive prefix to update the models. This paper extends their study by implementing hotfixing to mitigate another important type of undesired behavior: LLM4Code can complete many known buggy code; we aim at making LLM4Code not complete buggy code but complete the corresponding fixed code. First, we replicate the experiments of Jesse et al. (Jesse et al., 2023), confirming that models from a popular family (CodeGen) can memorize and produce many bugs, which is treated as undesired behaviors that should be mitigated. These known bugs have been fixed in open-source software; thus we hotfix the models and steer them to generate the corresponding fixed versions. To achieve this goal, we construct a dataset consisting of pairs of two code snippets: one containing bugs and the other with corresponding bugs being fixed. Then, we compute the diff (to get the deleted buggy code and the added fixed code) of each pair.
@@
We experiment a set of parameter-efficient fine-tuning (PEFT) methods to hotfix the models, which are designed to update the model with a small number of parameters, resulting in efficiency gains and making the deployment easier. We investigate a family of popular LLM4Code, CodeGen (Nijkamp et al., 2023), which is widely adopted in multiple studies (Jesse et al., 2023; Yang et al., 2024b; Li et al., 2023; Allal et al., 2023). By conducting experiments on models of two sizes (CodeGen-350M and CodeGen-2B), we gain the following insights about hotfixing LLM4Code. The optimal hotfixing setting (combining all three objectives and using LoRA for fine-tuning) can increase the generation of fixed code by up to 108.42% and decrease the generation of buggy code by up to 50.47%, which is more effective than the mitigation strategy adopted by Jesse et al. (Jesse et al., 2023). Although hotfixing reduces the model’s functional correctness on the HumanEval benchmark, statistical tests show that the difference is not statistically significant ($p$-values are greater than 0.05).
@@
  • Effectiveness: We implement a hotfix strategy by extending the work by He and Vechev (He and Vechev, 2023) with more PEFT methods on a new task, showing that it can effectively mitigate undesired behaviors with no significant impact on the functional performance.
@@ 3.3. Learning Objectives
However, in the context of hotfixing LLM4Code, directly training models on the entire updated examples may not be the appropriate option. The reason is highlighted in the work by He and Vechev (He and Vechev, 2023). In hotfixing, not each token in the entire updated examples is equally important, and the model should pay more attention to the added fixed code. Besides, the model should also learn not to generate the original buggy code and the model should cause minimal impact on the outputs that are not relevant to the bug-fixing updates. We explain the three objectives and the corresponding loss function design as follows.
@@ 3.3.2. Learning to Avoid Undesired Code
In the above equation, $w_{t}^{-}$ is the weight in computing the loss on token $x_{t}$, where $w_{t}^{-}=1$ if $x_{t}$ is the desired code and $w_{t}^{-}=0$ otherwise. Unlike the desired code, we want to increase the model loss on the undesired code to avoid them from being generated. However, simply maximizing the loss on the undesired code is ineffective. Our experiment shows that maximizing this loss will make the model lose its ability in generating meaningful code (the hotfixed models generate zero buggy code and zero fixed code), which is not desired in hotfixing. Instead of directly maximizing this loss, He and Vechev (He and Vechev, 2023) propose to adopt the loss function $\frac{L_{guided}}{L_{guided}+L_{unlearn}}$, which guides the optimizer to find a balance between minimizing the loss on desired code $L_{guided}$ and maximizing the loss on undesirable code $L_{unlearn}$. Ideally, to minimize this loss, the optimizer works to reduce $L_{guided}$ which encourages fixed code generation and to increase $L_{unlearn}$ which discourages the generation of buggy code. To make the model learn the desired code and avoid the undesired code, we define the Dual loss as follows:
@@ 3.3.3. Learning to Retrain Knowledge of Other Code
The fact that tuning language models on a new dataset can lead to performance degradation on other aspects has been observed in previous studies. For example, Liu et al. (Liu et al., 2016) point out that optimizing models on a new dataset can improve the average performance (e.g., BLEU score) but forgo coherence and fluency. This degeneration is often diagnosed as an effect of deviating too much from the original pre-trained model during optimization. Consequently, we want the model to behave similarly to the original model on the code that is not relevant to the undesired behavior. This is similar to the target of knowledge distillation (Tang et al., 2024; Shi et al., 2024), which aims to train a student model to mimic the behavior of the teacher model. In knowledge distillation, Kullback-Leibler (KL) divergence (Kullback and Leibler, 1951) is widely used to measure the behavioral difference between the two models and the student model learns to minimize the difference (Shi et al., 2023; Kim and Rush, 2016; Kim et al., 2024). He and Vechev also use it to mitigate the potential functional correctness decrease. The KL divergence loss is defined as follows:
@@ 7. Related Work
This is an extended replication study of the work by He and Vechev, which use contrastive prefix-tuning (Qian et al., 2022) to control the security-level of code generators. We extend their work on a new hotfixing task (i.e., reduce the number of generated buggy code and increase the number of fixed code) with more parameter-effect fine-tuning methods. We discuss the related works of other undesired behaviors in LLM4Code and efficient-tuning methods for models of code as follows.
@@ 7.2. Efficient-tuning for Models of Code
Weyssow et al. (Weyssow et al., 2024) use 5 parameter-efficient methods to fine-tune models for code generation tasks. We reuse their replication package to implement the first hotixing for mitigating undesired behavior. Prefix-tuning falls into the category of efficient-tuning methods for pre-trained models, which aim to utilize computationally-friendly techniques to efficiently adapt a model to a new dataset or task. Typical methods include in-context learning, prefix-tuning, adapter-tuning. Here we briefly introduce the related work about applying these methods to models of code. In-context learning (ICL) presents a small number of examples in the prompt to guide language models to conduct specific tasks. Gao et al. (Gao et al., 2023) evaluate ICL on code summarization, bug fixing, and program synthesis. They show that strategically designed context can lead to substantial improvements over widely-used demonstration construction methods. Huang et al. (Huang et al., 2023a) show that ICL can mitigate the social bias of code generated by language models. Wang et al. (Wang et al., 2022) apply prompt-tuning on three code intelligence tasks and show that prompt-tuning consistently outperforms fine-tuning in three investigated tasks. Wang et al. (Wang et al., 2023) evaluate adapter tuning for code search and summarization, showing that adapter tuning significantly outperforms full-model fine-tuning and effectively overcomes catastrophic forgetting. Choi and Lee (Choi and Lee, 2023) propose CodePrompt, using task-agnostic prefix tuning for program and language generation.
@@ 8. Conclusion and Future Work
In this study, we explore to what extent parameter-efficient fine-tuning (PEFT) techniques can be used to hotfix LLM4Code. We extend the work by He and Vechev (Qian et al., 2022) by evaluating four different PEFT techniques on the task of reducing the number of generated buggy code and increasing the number of fixed code. We conduct our experiments on the models in the CodeGen family. The empirical results demonstrate the effectiveness of hotfixing: the best combination leads to a significant increase in the generation of fixed code (up to 108.42%) and a notable reduction in the production of buggy code (up to 50.47%). We also conduct statistical analysis to confirm that hotfixing does not adversely impact the functional performance of the models on the HumanEval benchmark. We discuss the potential applications of the proposed method in various tasks, as well as its advantage in efficient deployment.